What is the equation of the line that contains the points with (x,y) coordinates (–3,7) and (5,–1) ?
Found 2 solutions by solver91311, richwmiller:
Answer by solver91311(24713)



Use the two-point form of the equation of a line:



Where and are the coordinates of the given points.

This will give you "an" equation of the line. You cannot determine "the" equation of a line because there are an infinite number of equivalent representations of any given line.

John


Answer by richwmiller(17219)

-1-7=(5-(-3))*m
-8=8m
-8/8=m
-1=m
y-7=-1(x-(-3)
y-7=-1x-4*3
y-7=-x-12
y=-x-5
This lines slant down from left to right.
when x=0 y=-5
when y=0 x=-5
BTW There are an infinite number of lines with those x's and y's not just one.
So the question should have been "What is an equation of a line etc?"